# A fast particle-field method for 3D Chemotaxis systemsJingyuan Hu, Zhongjian Wang, Jack Xin and Zhiwen Zhang

### v2

@@ 4.2 Relaxation of the Smoothness Assumption
In this section, we provide some conclusions indicating that under specific parameter combinations, the smoothness of $\hat{c}$ in Assumption 4.5 can be naturally guaranteed by the algorithm.
@@
Assume the particle count satisfies $P\geq H^{\frac{40}{13}}$ and the cutoff threshold is set as $H_{0}=H^{\frac{8}{13}}$. Then the condition (58) in Assumption 4.5 can be relaxed to:
@@
for any multi-index $\mathbf{k}$ such that $|\mathbf{k}|=\eta$ where $\eta=1,2,3$, and some constants $M_{\eta},Q_{\eta},R_{\eta}$ that are independent of $H$, $H_{0}$, and $P$.
@@
We show that $\|D_{\mathbf{k}}\hat{c}\|_{2}$ is controlled by $\|D_{\mathbf{k}}c\|_{2}$. For $\eta=1,2,3$, by the Parseval theorem,
@@
By Lemmas 4.10 and 4.13, we have the estimate
@@
$\leq\left(\frac{2\pi}{L}\right)^{2\eta-2}\left(C_{4}H_{0}^{2\eta+7}H^{-8}+C_{5}H_{0}^{2\eta-1}P^{-1}+C_{6}H_{0}^{2\eta-6}+C_{7}(D^{(n)})^{2}H_{0}^{2\eta-2}\right)$
where the constants $C_{4}=M_{0}^{2}N_{2}^{-1}C_{1}^{2}L^{-1},C_{5}=M_{0}^{2}N_{2}^{-1}C_{2}^{2}L^{-1},C_{6}=L^{4}(2\pi)^{-4}N_{1,2},\\ C_{7}=2M_{0}^{2}T$ are independent of $H,H_{0},P,M_{\eta}$. For the parameters $P\geq H^{\frac{40}{13}}$ and $H_{0}=H^{\frac{8}{13}}$, we have $\|D_{\mathbf{k}}\hat{c}^{(n)}\|_{2}\leq 2\|D_{\mathbf{k}}c^{(n)}\|_{2}+B_{\eta}$ while $D^{(n)}\leq H^{-\frac{16}{13}}$, for some $B_{\eta}$ independent of $H,H_{0},P,M_{\eta}$ where $\eta=1,2,3$.
@@
Assume the particle count satisfies $P\geq H^{\frac{64}{13}}$ and the cutoff threshold is set as $H_{0}=H^{\frac{8}{13}}$. For the timestep converging to zero more slowly, specifically, $H_{\tau}:=\tau^{-\frac{1-\sigma}{2}}<H_{0}$ for some $\sigma>0$, the condition (58) in Assumption 4.5 can be relaxed to:
@@
for any multi-index $\mathbf{k}$ such that $|\mathbf{k}|=\eta$ where $\eta=1,2,3$.
@@
We verify that the conditions of Lemma 4.19 are satisfied. Let $\hat{c}^{(n)}_{\text{low}}=\sum_{\|\mathbf{q}\|\leq H_{\tau}}\hat{\alpha}_{\mathbf{q}}^{(n)}e^{-i\frac{2\pi}{L}\mathbf{q}\cdot\mathbf{x}}$, $\hat{c}^{(n)}_{\text{high}}=\sum_{H_{\tau}<\|\mathbf{q}\|\leq H_{0}}\hat{\alpha}_{\mathbf{q}}^{(n)}e^{-i\frac{2\pi}{L}\mathbf{q}\cdot\mathbf{x}}$. Then $\|D_{\mathbf{k}}\hat{c}^{(n)}\|_{\infty}\leq\\ \|D_{\mathbf{k}}\hat{c}^{(n)}_{\text{low}}\|_{\infty}+\|D_{\mathbf{k}}\hat{c}^{(n)}_{\text{high}}\|_{\infty}$. For the low-frequency part, by the Parseval Theorem,
@@
For the high-frequency part, we verify that $\left|\hat{\alpha}_{\mathbf{q}}^{(n)}\right|$ ($H_{\tau}<\|\mathbf{q}\|\leq H_{0}$) is limited by the diffusion $\mathbf{W}_{p}^{(n)}\sim\mathcal{N}(0,2\mu\tau I_{3})$ at every step. Recall that
@@
Let the summation term of the $p$-th particle $\hat{\beta}_{\mathbf{q},p}^{(n)}=e^{-i\mathbf{X}_{p}^{(n)}\cdot\frac{2\pi}{L}\mathbf{q}}$. Then $\left|\mathbb{E}(\hat{\beta}_{\mathbf{q},p}^{(n)})\right|=e^{-\mu\tau\left(\frac{2\pi}{L}\mathbf{q}\right)^{2}}$, $\text{Var}(\hat{\beta}_{\mathbf{q},p}^{(n)})=1-e^{-\mu\tau\left(\frac{2\pi}{L}\mathbf{q}\right)^{2}}\leq 1$. Since $\hat{\beta}_{\mathbf{q},p}^{(n)}$ are independent and $\mathbb{E}(\left|\hat{\beta}_{\mathbf{q},p}^{(n)}\right|^{\xi})\\ =1$ for every $\xi>0$, from Lindeberg’s condition, we have $\left|\hat{\beta}_{\mathbf{q}}^{(n)}\right|\leq\frac{M_{0}}{L^{3}}(e^{-\mu\tau\left(\frac{2\pi}{L}\mathbf{q}\right)^{2}}+\frac{S_{1}}{\sqrt{P}})$ with probability $1-e^{-\frac{S_{1}^{2}}{2}}$. Assume that this bound holds for all $q\in U_{0}\setminus U_{\tau}$ and for $1\leq n\leq N_{T}$, with probability at least $1-H_{0}^{3}N_{T}e^{-\frac{S_{1}^{2}}{2}}$. Similar to (88), the update of $\hat{\alpha}_{\mathbf{q}}$ satisfies
@@
where $b^{(n-1)}_{\mathbf{q}}\leq\frac{M_{0}}{L^{3}}CH^{-4}$ is the interpolation error. Similar to (81), this accumulates into
@@
Summing over all coefficients, we obtain:
@@
Since $|\mathbf{q}|^{2}\tau>|\mathbf{q}|^{\frac{2\sigma}{1-\sigma}}$ for $|\mathbf{q}|>H_{\tau}$, the first term $\sum_{\|\mathbf{q}\|>H_{\tau}}\left|\mathbf{q}\frac{2\pi}{L}\right|^{\eta-2}\frac{M_{0}}{L^{3}}e^{-\mu\tau\left(\frac{2\pi}{L}\mathbf{q}\right)^{2}}\leq\frac{M_{0}}{L^{3}}C_{\sigma}$ for some $C_{\sigma}>0$, as it tends to zero while $H_{\tau}\to\infty$. Thus
@@
From (108) and (113), with the conditions $H_{0}=H^{\frac{8}{13}}$ and $P\geq H_{0}^{8}$, we obtain the conclusion to be proved.
@@
The presented error analysis critically relies on the non-blow-up assumptions embedded in the framework. In scenarios where solutions develop singularities or lose regularity (e.g., near a delta-type blow-up), the square-integrability condition and the smoothness requirement in Assumption 4.5 fail to hold.
@@ 5 Numerical Experiments
This section presents numerical validation and application of the SIPF-PIC framework. We first establish convergence properties through systematic analysis of particle-grid resolution dependencies, followed by
demonstrations of the framework’s capacity to resolve finite-time blow-up dynamics.
The section is organized as follows. Subsection 5.1 quantifies algorithmic convergence through error analysis relative to particle count $P$ and grid resolution $H$ per dimension, and subsequently compares computational performance using identical Keller-Segel parameters. Building on these foundations, Subsection 5.2 verifies the theoretical threshold $M_{\text{critical}}=8\pi$ in 2D elliptic systems and identifies an empirical threshold in 3D parabolic systems. We then explore blow-up dynamics: Subsection 5.3 reveals three mass-dependent phases in tetrahedral symmetric systems—subcritical dispersion, intermediate unified blowup, and supercritical multi-stage singularity formation. Finally, Subsection 5.4 demonstrates radial concentration into singular ring structures with post-collapse instabilities.
5.1 Convergence of SIPF-PIC with Respect to Discretization
@@
with total mass $M_{0}=80$. The model parameters in (1) are set as $\mu=\chi=1$, $\epsilon=10^{-4}$, $k=10^{-1}$, with final time $T=2\times 10^{-2}$ and timestep $\tau=10^{-5}$, total steps $N_{T}=2000$.
By exploiting radial symmetry, we can reduce the three-dimensional system to a one-dimensional radial coordinate $r$ through the transformation
@@
(116) $\begin{cases}\displaystyle\rho_{t}=\frac{2\mu}{r}\rho_{r}+\mu\rho_{rr}-\frac{2\chi}{r}\rho c_{r}-\chi\rho_{r}c_{r}-\chi\rho c_{rr},\\ \displaystyle\epsilon c_{t}=\frac{2}{r}c_{r}+c_{rr}-k^{2}c+\rho.\end{cases}$
with boundary conditions $\rho_{r}|_{r=0}=c_{r}|_{r=0}=0$.
While the problem defined in (114) is posed in $\mathbb{R}^{3}$, the particle dynamics remain localized within $|\mathbf{x}|<2$ throughout the simulation. The periodic setting in $\Omega=[-10,10]^{3}$ (i.e., $L=20$) thus suffices to approximate the unbounded problem, as there is almost no density distribution outside of $\Omega$.
@@
Convergence studies for SIPF-PIC are conducted by varying $P$ and $H$ while comparing with this reference. We use the cumulative distribution function
@@
at the final time $T=2\times 10^{-2}$ as the comparative criterion. Since $m(0,T;\rho)=0$ and $m(\infty,T;\rho)=1$ for every mass-preserving density function $\rho$, we compute the $W^{1}$ distance between the results of SIPF-PIC and the reference solution as the error. Specifically, for the density $\hat{\rho}$ defined by (6), we have $m(r,T,\hat{\rho})=\frac{1}{P}\sum_{p=1}^{P}\mathbf{1}_{\{|X_{p}^{(N_{T})}|\leq r\}},$ and for a radially symmetric density $\rho(\mathbf{x},t)=\psi(|\mathbf{x}|,t)$, we have $m(r,T,\rho)=\frac{4\pi}{M_{0}}\int_{0}^{r}\psi(s,t)s^{2}\,ds.$ To compute the Wasserstein distance, we obtain the quantiles $q_{\frac{j}{n}}$ and $\hat{q}_{\frac{j}{n}}$ for $n=10^{4}$ and $j=0,1,\ldots,n-1$, limited to the interval $[0,R]$, where $R=50$ is a sufficiently large radius. Specifically, we have $m(q^{*}_{\frac{j}{n}},T,\rho)=\frac{j}{n},\quad m(\hat{q}^{*}_{\frac{j}{n}},T,\hat{\rho})=\frac{j}{n}$ with $q_{\frac{j}{n}}=\min(q^{*}_{\frac{j}{n}},R),\quad\hat{q}_{\frac{j}{n}}=\min(\hat{q}^{*}_{\frac{j}{n}},R).$
Finally, we define the Wasserstein $W^{1}$ error:
@@
We systematically tested all combinations of $P\in\{2^{11},\ldots,2^{20}\}$ and $H\in\{8,16,\\ 32,64,128,256\}$. In these examples, we set $H_{0}=H$, meaning that no frequency cutoff is applied.
Figure 3 demonstrates the convergence behavior using fourth-order Particle-to-Grid and second-order Grid-to-Particle interpolations, quantifying convergence rates with respect to individual parameters. Two complementary studies isolate parametric dependencies: Fixing the grid resolution at $H_{\max}$ yields particle-count convergence rates via the slope of $\log_{2}\mathcal{E}$ versus $\log_{2}P$, measured as 0.4207-0.4207, which closely matches the theoretical order of $1/2$ derived in Subsection 4; while fixing $P_{\max}$ determines spatial convergence orders through the slope of $\log_{2}\mathcal{E}$ versus $\log_{2}H$, measured as 1.4975-1.4975, consistent with the theoretical order of $16/13$.
(a) Error Convergence Against Particle Count ($P$) With Fixed Grid Resolution
(b) Error Convergence Against Grid Resolution ($H$) With Fixed Particle Count
Figure 3: Convergence characteristics with dual parameter analysis.
In addition to the convergence in Figure 3, we investigate a comparison between the original SIPF method and our new SIPF-PIC framework (using second or fourth order interpolation) through timing and error measurements for the Keller-Segel system, using identical parameter sets and initial conditions as in Section 5.1. The test cases employ grid resolutions $H\in\{8,24,64,256\}$ and particle counts $P\in\{2^{11},2^{14},2^{17},2^{20}\}$, with results tabulated in Table 1.
Table 1: Time-error profiles of SIPF and SIPF-PIC with different interpolation orders
@@
The complexity reduction from $O(PH^{3})$ to $O(P+H^{3}\log H)$ through SIPF-PIC’s localized particle-grid interpolations and spectral field solutions fundamentally decouples particle-field interactions, achieving accelerated parameter exploration while preserving numerical accuracy without reintroducing $PH^{3}$ bottlenecks. This complexity separation permits simulations at extreme resolutions ($H\geq 256,P\geq 2^{20}$), with controlled error growth, overcoming prior SIPF limitations through adaptive accuracy-efficiency tradeoffs.
We also demonstrate that the algorithm is consistent with respect to $H$. When the number of particles $P$ is fixed and the grid resolution $H$ increases, the number of grid points $H^{3}$ may become significantly larger than $P$. In this scenario, according to Theorem 4.16, our algorithm still yields convergent results. We fix $P=10^{4}$ as in the original SIPF algorithm and set the truncation frequency $H_{0}=16$, while increasing $H$ from $16$ to $256$. The error of each configuration compared to the reference solution is shown in Fig. 4. As $H$ increases, the error converges to a value determined by $P$ and $H_{0}$.
Figure 4: Convergence with increasing grid resolution $H$ under a fixed particle count ($P=10^{4}$) and truncation frequency $H_{0}=16$.

### v3

@@ 4.2 Relaxation of the smoothness assumption
In this section, we establish several conclusions showing that under specific parameter combinations, the smoothness of $\hat{c}$ in Assumption 4.5 can be naturally guaranteed by the algorithm itself.
@@
Assume the particle number satisfies $P\geq H^{\frac{40}{13}}$ and the frequency cutoff threshold is set to $H_{0}=H^{\frac{8}{13}}$. Then, the condition (50) in Assumption 4.5 can be relaxed to:
@@
for any multi-index $\mathbf{k}$ with $|\mathbf{k}|=\eta$, where $\eta=1,2,3$, and for some constants $M_{\eta},Q_{\eta},R_{\eta}$ that are independent of $H$, $H_{0}$, and $P$.
@@
We show that $\|D_{\mathbf{k}}\hat{c}\|_{2}$ is controlled by $\|D_{\mathbf{k}}c\|_{2}$. For $\eta=1,2,3$, by Parseval’s theorem,
@@
By Lemmas 4.10 and 4.13, we can obtain the estimate
@@
$\leq\left(\frac{2\pi}{L}\right)^{2\eta-2}\left(C_{4}H_{0}^{2\eta+7}H^{-8}+C_{5}H_{0}^{2\eta-1}P^{-1}+C_{6}H_{0}^{2\eta-6}+C_{7}(D^{(n)})^{2}H_{0}^{2\eta-2}\right),$
where the constants $C_{4}=M_{0}^{2}N_{2}^{-1}C_{1}^{2}L^{-1}$, $C_{5}=M_{0}^{2}N_{2}^{-1}C_{2}^{2}L^{-1}$, $C_{6}=L^{4}(2\pi)^{-4}N_{1,2}$, and $C_{7}=2M_{0}^{2}T$ are independent of $H$, $H_{0}$, $P$, and $M_{\eta}$.
For the parameters $P\geq H^{\frac{40}{13}}$ and $H_{0}=H^{\frac{8}{13}}$, we have $\|D_{\mathbf{k}}\hat{c}^{(n)}\|_{2}\leq 2\|D_{\mathbf{k}}c^{(n)}\|_{2}+B_{\eta}$ for $D^{(n)}\leq H^{-\frac{16}{13}}$, for some $B_{\eta}$ independent of $H$, $H_{0}$, $P$, and $M_{\eta}$ with $\eta=1,2,3$.
@@
Assume the particle number satisfies $P\geq H^{\frac{64}{13}}$ and the frequency cutoff threshold is set to $H_{0}=H^{\frac{8}{13}}$. For a timestep converging to zero more slowly, specifically, $H_{\tau}:=\tau^{-\frac{1-\sigma}{2}}<H_{0}$ for some $\sigma>0$, the condition (50) in Assumption 4.5 can be relaxed to
@@
for any multi-index $\mathbf{k}$ with $|\mathbf{k}|=\eta$, where $\eta=1,2,3$.
@@
We verify that the conditions of Lemma 4.19 are satisfied. Let $\hat{c}^{(n)}_{\text{low}}=\sum_{\|\mathbf{q}\|\leq H_{\tau}}\hat{\alpha}_{\mathbf{q}}^{(n)}e^{-i\frac{2\pi}{L}\mathbf{q}\cdot\mathbf{x}}$ and $\hat{c}^{(n)}_{\text{high}}=\sum_{H_{\tau}<\|\mathbf{q}\|\leq H_{0}}\hat{\alpha}_{\mathbf{q}}^{(n)}e^{-i\frac{2\pi}{L}\mathbf{q}\cdot\mathbf{x}}$. Then, $\|D_{\mathbf{k}}\hat{c}^{(n)}\|_{\infty}\leq\\ \|D_{\mathbf{k}}\hat{c}^{(n)}_{\text{low}}\|_{\infty}+\|D_{\mathbf{k}}\hat{c}^{(n)}_{\text{high}}\|_{\infty}$. For the low-frequency part, by the Parseval Theorem,
@@
For the high-frequency part, we verify that $\left|\hat{\alpha}_{\mathbf{q}}^{(n)}\right|$ (for $H_{\tau}<\|\mathbf{q}\|\leq H_{0}$) is bounded by the diffusion term $\mathbf{W}_{p}^{(n)}\sim\mathcal{N}(0,2\mu\tau I_{3})$ at each step. Recall that
@@
Let the summation term for the $p$-th particle be $\hat{\beta}_{\mathbf{q},p}^{(n)}=e^{-i\mathbf{X}_{p}^{(n)}\cdot\frac{2\pi}{L}\mathbf{q}}$. Then, we can obtain that $\left|\mathbb{E}(\hat{\beta}_{\mathbf{q},p}^{(n)})\right|=e^{-\mu\tau\left(\frac{2\pi}{L}\mathbf{q}\right)^{2}}$, and $\text{Var}(\hat{\beta}_{\mathbf{q},p}^{(n)})=1-e^{-\mu\tau\left(\frac{2\pi}{L}\mathbf{q}\right)^{2}}\leq 1$. Since $\hat{\beta}_{\mathbf{q},p}^{(n)}$ are independent and $\mathbb{E}(\left|\hat{\beta}_{\mathbf{q},p}^{(n)}\right|^{\xi})=1$ for all $\xi>0$, by Lindeberg’s condition, we have $\left|\hat{\beta}_{\mathbf{q}}^{(n)}\right|\leq\frac{M_{0}}{L^{3}}(e^{-\mu\tau\left(\frac{2\pi}{L}\mathbf{q}\right)^{2}}+\frac{S_{1}}{\sqrt{P}})$ with probability $1-e^{-\frac{S_{1}^{2}}{2}}$.
Assume that this bound holds for all $q\in U_{0}\setminus U_{\tau}$ and for $1\leq n\leq N_{T}$, with probability at least $1-H_{0}^{3}N_{T}e^{-\frac{S_{1}^{2}}{2}}$. Similar to (80), the update rule for $\hat{\alpha}_{\mathbf{q}}$ satisfies
@@
where $b^{(n-1)}_{\mathbf{q}}\leq\frac{M_{0}}{L^{3}}CH^{-4}$ is the interpolation error. Similar to (73), this accumulates to
@@
Summing over all coefficients, we obtain
@@
Since $|\mathbf{q}|^{2}\tau>|\mathbf{q}|^{\frac{2\sigma}{1-\sigma}}$ for $|\mathbf{q}|>H_{\tau}$, the first term $\sum_{\|\mathbf{q}\|>H_{\tau}}\left|\mathbf{q}\frac{2\pi}{L}\right|^{\eta-2}\frac{M_{0}}{L^{3}}e^{-\mu\tau\left(\frac{2\pi}{L}\mathbf{q}\right)^{2}}\\ \leq\frac{M_{0}}{L^{3}}C_{\sigma}$ for some constant $C_{\sigma}>0$, as it tends to zero as $H_{\tau}\to\infty$. Thus,
@@
From (100) and (105), under the conditions $H_{0}=H^{\frac{8}{13}}$ and $P\geq H_{0}^{8}$, we conclude the proof.
@@
The presented error analysis critically relies on the non-blow-up assumptions embedded in the framework. In scenarios where solutions develop singularities or lose regularity (e.g., near a delta-type blow-up), the square-integrability condition and the smoothness requirement in Assumption 4.5 no longer hold.
@@ 5 Numerical experiments
This section presents numerical results to validate the convergence analysis and demonstrate the performance of the SIPF-PIC method. We first establish convergence properties through systematic analysis of particle-grid resolution dependencies, followed by demonstrations of the framework’s capacity to resolve finite-time blow-up dynamics.
The section is organized as follows. Subsection 5.1 quantifies algorithmic convergence through error analysis with respect to the particle number $P$ and grid resolution $H$ per dimension, and subsequently compares computational performance using identical Keller-Segel parameters. Building on these foundations, Subsection 5.2 verifies the theoretical threshold $M_{\text{critical}}=8\pi$ in 2D elliptic systems and identifies an empirical threshold in 3D parabolic systems. We then explore blow-up dynamics: Subsection 5.3 reveals three mass-dependent phases in tetrahedral symmetric systems, including subcritical dispersion, intermediate unified blowup, and supercritical multistage singularity formation. Finally, Subsection 5.4 demonstrates radial concentration into singular ring structures that exhibit post-collapse instabilities.
5.1 Convergence of the SIPF-PIC method with respect to discretization
@@
with total mass $M_{0}=80$. The model parameters in (1) are set as $\mu=\chi=1$, $\epsilon=10^{-4}$, $k=10^{-1}$, with final time $T=2\times 10^{-2}$, timestep $\tau=10^{-5}$, and total time steps $N_{T}=2000$.
By exploiting radial symmetry, we can reduce the 3D system to a 1D radial coordinate $r$ through the following transformation
@@
(108) $\begin{cases}\displaystyle\rho_{t}=\frac{2\mu}{r}\rho_{r}+\mu\rho_{rr}-\frac{2\chi}{r}\rho c_{r}-\chi\rho_{r}c_{r}-\chi\rho c_{rr},\\ \displaystyle\epsilon c_{t}=\frac{2}{r}c_{r}+c_{rr}-k^{2}c+\rho,\end{cases}$
together with boundary conditions $\rho_{r}|_{r=0}=c_{r}|_{r=0}=0$.
While the problem defined in (106) is posed in $\mathbb{R}^{3}$, the particle dynamics remain localized within $|\mathbf{x}|<2$ during the entire simulation. The periodic setting in $\Omega=[-10,10]^{3}$ (i.e., $L=20$) thus suffices to approximate the unbounded problem, as there is almost no density distribution outside of the domain $\Omega$.
@@
Convergence studies for the SIPF-PIC method are conducted by varying $P$ and $H$ while comparing the numerical results with this reference solution. We use the cumulative distribution function
@@
at the final time $T=2\times 10^{-2}$ as the comparative criterion. Since $m(0,T;\rho)=0$ and $m(\infty,T;\rho)=1$ for every mass-preserving density function $\rho$, we compute the $W^{1}$ distance between the results obtained by the SIPF-PIC method and the reference solution as the error metric. Specifically, for the density $\hat{\rho}$ defined by (6), we have $m(r,T,\hat{\rho})=\frac{1}{P}\sum_{p=1}^{P}\mathbf{1}_{\{|X_{p}^{(N_{T})}|\leq r\}},$ and for a radially symmetric density $\rho(\mathbf{x},t)=\psi(|\mathbf{x}|,t)$, we have $m(r,T,\rho)=\frac{4\pi}{M_{0}}\int_{0}^{r}\psi(s,t)s^{2}\,ds.$ To compute the Wasserstein distance, we obtain the quantiles $q_{\frac{j}{n}}$ and $\hat{q}_{\frac{j}{n}}$ for $n=10^{4}$ and $j=0,1,\ldots,n-1$, restricted to the interval $[0,R]$, where $R=50$ is a sufficiently large radius. Specifically, we have $m(q^{*}_{\frac{j}{n}},T,\rho)=\frac{j}{n}$ and $m(\hat{q}^{*}_{\frac{j}{n}},T,\hat{\rho})=\frac{j}{n}$ with $q_{\frac{j}{n}}=\min(q^{*}_{\frac{j}{n}},R)$ and $\hat{q}_{\frac{j}{n}}=\min(\hat{q}^{*}_{\frac{j}{n}},R)$.
Finally, we define the Wasserstein $W^{1}$ error as
@@
We test all combinations of $P\in\{2^{11},\ldots,2^{20}\}$ and $H\in\{8,16,32,64,128,256\}$. In these simulations, we set $H_{0}=H$, meaning that no frequency cutoff is applied.
Figure 3 demonstrates the convergence behavior of the SIPF-PIC method using fourth-order particle-to-grid and second-order grid-to-particle interpolations, quantifying convergence rates with respect to the individual parameters. Two complementary studies isolate the parametric dependencies: fixing the grid resolution at $H_{\max}$ yields the particle number convergence rate via the slope of $\log_{2}\mathcal{E}$ versus $\log_{2}P$, with a measured value of 0.4207-0.4207, which closely matches the theoretical order of $1/2$ derived in Subsection 4; fixing $P_{\max}$ determines the spatial convergence order through the slope of $\log_{2}\mathcal{E}$ versus $\log_{2}H$, with a measured value of 1.4975-1.4975, which closely matches the theoretical order of $16/13$.
(a) Error convergence against particle number ($P$) with fixed grid resolution.
(b) Error convergence against grid resolution ($H$) with fixed particle number.
Figure 3: Convergence characteristics under the dual parameters of particle number and grid resolution.
In addition to the convergence results presented in Figure 3, we investigate a performance comparison between the original SIPF method and our new SIPF-PIC method (using second- or fourth-order interpolation) through timing and error measurements for the KS system, using identical parameter sets and initial conditions to those in Section 5.1. The test cases employ grid resolutions $H\in\{8,24,64,256\}$ and particle numbers $P\in\{2^{11},2^{14},2^{17},2^{20}\}$, with results tabulated in Table 1.
Table 1: Time-error profiles of SIPF and SIPF-PIC with different interpolation orders.
@@
The complexity reduction from $O(PH^{3})$ to $O(P+H^{3}\log H)$ through the SIPF-PIC method’s localized particle-grid interpolations and spectral field solutions fundamentally decouples particle-field interactions, achieving accelerated parameter exploration while preserving numerical accuracy and avoiding the reintroduction of $PH^{3}$ bottlenecks. This complexity separation permits high-resolution simulations at extreme resolutions (e.g., $H\geq 256$ and $P\geq 2^{20}$) with controlled error growth, overcoming prior SIPF limitations via adaptive accuracy-efficiency tradeoffs.
We also demonstrate that the algorithm is consistent with respect to $H$. When the number of particles $P$ is fixed and the grid resolution $H$ is increased, the number of grid points $H^{3}$ may become significantly larger than $P$. In this scenario, according to Theorem 4.16, our algorithm still yields convergent results. We fix $P=10^{4}$ (consistent with the original SIPF method) and set the truncation frequency $H_{0}=16$, while increasing $H$ from $16$ to $256$. The error of each configuration compared to the reference solution is shown in Fig. 4. As $H$ increases, the error converges to a constant value determined by $P$ and $H_{0}$.
Figure 4: Convergence with increasing grid resolution $H$ under a fixed particle number ($P=10^{4}$) and truncation frequency $H_{0}=16$.